OpenStudy (anonymous):

solve for x. show all steps leading to your solution.and check your solution 34x-18=10x-9

jimthompson5910 (jim_thompson5910):

34x-18=10x-9 34x=10x-9+18 34x-10x=-9+18 24x=-9+18 24x=9 x=9/(24) x=3/8

OpenStudy (anonymous):

34x-18=10x-9 24x=9 x= 9/24 x= 3/8